# Design of a Wearable Parallel Electrical Impedance Imaging System for Healthcare

### v1

@@ Abstract
A wireless wearable electrical impedance tomography (EIT) system has been developed using the AD5933 for real-time lung respiration imaging. The system uses a current injection method tailored to the human body’s impedance characteristics. It applies a voltage excitation and measures the resulting current as the voltage passes through the body. Additionally, measures are taken to suppress the effects of parasitic capacitance, which can lead to signal oscillations and leakage currents. Additionally, to improve data acquisition speed, five AD5933 units are used for parallel measurements, with multiple measures taken to ensure high synchronization during the parallel acquisition process. The results demonstrate conductivity images generated from the EIT system, with data collected from a water tank, human lungs during respiration, and a human calf at rest, confirming that this portable EIT system can measure biological tissues with high precision and low cost.
@@ 1 Introduction
EIT is widely utilized in clinical fields such as lung function monitoring [3, 4], brain function monitoring [5], cancer detection [6], and human-machine interfaces [7]. Lung function monitoring is particularly prevalent, as EIT provides real-time visualizations of pulmonary gas distribution, assisting clinicians in optimizing ventilator settings, thus reducing risks associated with ventilator-induced lung injury.
@@
Commercial EIT systems include Sentec LuMon [8], Dräger Pulmovista 500 [9], Timpel Enlight 2100 [10], Gense [11], and Sciospec’s EIT16 and EIT32/64/128+ systems [12]. Notably, the Dräger Pulmovista 500 system offers real-time visualization of pulmonary gas distribution, helping clinicians assess lung collapse, overinflation, and ventilation heterogeneity. Sciospec’s systems typically employ 16 electrodes but also support configurations up to 256 electrodes, achieving real-time frame rates of 100 fps with frequencies ranging from 100 Hz to 1 MHz.
@@
Recent advancements in EIT technology include systems like Sheffield Mk3.5 [13], ACT-4 [14], ACT-5 [15], OXBACT-3 [16], OXBACT-5 [17], UCLH Mk2 [18], UCLH Mk2.5 [19], KHU Mark2 [20], KHU Mark2.5 [21, 22], mfEIT [23], and Dartmouth College-developed systems [24, 25, 26]. Among these, the Sheffield Mk3.5 system can measure impedance at 30 frequency points ranging from 2 kHz to 1.6 MHz, with an average signal-to-noise ratio (SNR) of 40 dB across all frequencies. This system utilizes 8 electrodes and includes 8 identical data acquisition boards, each equipped with a digital signal processor (DSP) to generate driving frequencies and perform fast Fourier transform (FFT) demodulation. The ACT-5 system, employing 32 electrodes, can display conductivity distribution images at a rate of 27 frames per second. Its adaptive current source can apply fully programmable current patterns over a frequency range of 5 kHz to 500 kHz. Additionally, it can simultaneously display electrocardiogram (ECG) readings during the imaging process. The KHU Mark2.5 system features automatic calibration, continuous operation at 250 kHz frequency for 24 hours, and maintains an SNR above 80 dB. This FPGA-based system includes multiple impedance measurement modules, each containing an independent constant current source, a differential voltage measurement module, and a current-source calibrator.
@@
Although the electrical impedance tomography (EIT) devices mentioned above offer excellent imaging performance, their high costs limit their widespread adoption. As a result, developing a low-cost, easy-to-implement EIT system is crucial. In clinical medicine, EIT is primarily used for real-time lung function monitoring. Therefore, the device must be capable of capturing respiratory changes in the lungs while also being compact, portable, and capable of fast data acquisition to meet the demands of dynamic imaging.
@@
This paper aims to develop a portable and low-cost EIT system based on the AD5933 chip. Firstly, according to the impedance characteristics of the human body, current injection is realized by applying an excitation voltage to the object under test and measuring the corresponding excitation current. Compared to traditional methods using a current source, this approach does not require the complete elimination of the DC offset inherent in current sources. Meanwhile, to tackle issues such as oscillations and leakage currents caused by parasitic capacitances in multiplexers, the system implements various optimization measures, enhancing stability at higher frequencies. Secondly, to improve data acquisition speed, the system employs five AD5933s operating in parallel. The TCA9548A is used to multiplex the I2C bus, while the ICS553 provides a unified external clock to all AD5933s, ensuring high synchronization during parallel acquisition. Finally, the PCB layout was optimized for a more compact design. Additionally, Bluetooth wireless communication and lithium battery power supply were integrated to further enhance the device’s portability. The structure of this paper is organized as follows: Section II introduces the developed EIT system and its specifications. Section III presents experimental results obtained from measurements on a water tank and a male subject. Section IV provides a brief summary.
@@ 2.1 System Architecture
The impedance hardware acquisition system is designed around the AD5933, which integrates a high-precision direct digital synthesizer (DDS) capable of generating sine wave signals within a frequency range of 1 kHz to 100 kHz. Additionally, the AD5933 incorporates a 12-bit analog-to-digital converter (ADC) for high-resolution analog signal acquisition. Most importantly, it includes a digital signal processor (DSP) that allows for accurate measurement of both the amplitude and phase of the sine wave signal.
@@
The overall architecture of the EIT system is shown in Fig.1(a), and the physical prototype in Fig.1(b), with all circuit modules labeled. The system consists of five AD5933 units, a control module, a power supply module, a Bluetooth transmission module, a voltage amplification module, an I-V conversion module, five voltage measurement modules, an I2C bus control module, an external clock module, and multiple multiplexer modules. During the impedance acquisition process, the STM32 first controls the first AD5933 via the I2C bus to generate a sine voltage signal, which is then amplified into a sine voltage excitation by the voltage amplification module. The output of the voltage amplification module and the current input of the I-V conversion circuit are connected to two electrodes for voltage excitation through two multiplexer modules. Meanwhile, the AD5933 selects either the output terminal connected to the boundary voltage measurement or the output terminal of the I-V conversion circuit to acquire either the boundary voltage data or the current data generated by the sine voltage excitation passing through the object under test. The STM32 controls the selection of different pins on the multiplexer chip, enabling a complete scan of the object under test. The AD5933 transmits the processed voltage and current data to the STM32 via the I2C bus. Finally, the STM32 sends the calculated impedance data to the computer software for image reconstruction via a serial bus or Bluetooth.
@@
The system offers high flexibility, allowing configuration into various measurement modes based on actual requirements. By adjusting the jumper settings, the enable (EN) pins of the multiplexers connected to the voltage detection circuitry can be set to either a high or low level, enabling or disabling multiplexer channels. This setup helps avoid signal conflicts and interference during measurements, while also reducing parasitic capacitance effects from unused multiplexers. It can operate in a two-terminal serial measurement mode for basic impedance detection, or switch to a four-terminal serial measurement mode, where current is injected through two electrodes and voltage is measured across the other two, thereby minimizing the effect of contact resistance on the results and enhancing accuracy. Additionally, the system supports a four-terminal semi-parallel measurement mode, enabling simultaneous voltage measurements across multiple electrodes, further accelerating the data acquisition process.
@@
In the two-terminal series mode of measurement, the impedance is calculated based on the excitation current generated by the fixed excitation voltage passing through the test object. At this moment, the system architecture is shown in Fig.2.
@@
In the four-terminal series measurement mode, a single voltage excitation can only obtain the voltage of one electrode or the current generated by the excitation voltage passes through the object. Therefore, channel switching via multiplexers is necessary to measure all boundary voltages and excitation currents in the adjacent excitation mode. At this point, the system architecture is shown in Fig.3.
@@
In the four-terminal semi-parallel measurement mode, by using five AD5933 chips in parallel, the system can simultaneously obtain the voltage of four electrodes and the current generated by the excitation voltage passing through the object during a single current injection. Through multiple channel switching, all boundary voltages and excitation currents under the adjacent excitation mode can be measured. At this point, the system architecture is shown in Fig.4.
Figure 4: EIT System Architecture in the Four-Terminal Semi-Parallel Mode.
@@ 2.2.1 Circuit structure analysis
An excitation voltage is applied to the test object, and the resulting excitation current is measured. In this manner, the measured current represents the actual excitation current injected into the test object. In contrast to the method that uses a fixed-amplitude current signal for excitation, this approach does not depend on the high output impedance of the current source to maintain current amplitude stability, nor does it require ensuring high linearity between the current amplitude and the control signal to guarantee measurement accuracy. Additionally, since a fixed-amplitude voltage excitation is used, the circuit does not need to dynamically adjust the output current based on the impedance characteristics of the test object. Another advantage of the voltage excitation method is that even if there is a small amount of DC bias in the excitation voltage, the resulting DC bias current will be minimal due to the human body’s extremely high impedance to DC. Therefore, there is no need to completely eliminate the DC bias voltage. Moreover, the circuit design is simple and reliable, requiring only a 3.3V power supply for stable operation.
@@
Fig.5 shows the voltage excitation circuit, and the output voltage of the circuit is:
@@
When $R_{2}=R_{1}$, the output signal $V_{0}(s)$ has the same AC amplitude as the input signal $V_{i}(s)$, and the DC bias of the output signal $V_{0}(s)$ is $\frac{Vcc}{2}$.
@@
Fig.6 shows the I-V conversion circuit, and the output voltage of the circuit is:
@@
The output terminal of the voltage excitation circuit and the current input terminal of the I-V conversion circuit are connected to both ends of the test object. The output signal of the voltage excitation circuit has an AC amplitude of 1.98 $V_{p-p}$ and a DC bias level of $\frac{Vcc}{2}$. Meanwhile, the voltage at the inverting input terminal of the operational amplifier in the I-V conversion circuit is also $\frac{Vcc}{2}$. Therefore, the DC voltage across the test object is 0V, while the AC voltage amplitude is 1.98 $V_{p-p}$.
@@ 2.2.2 Circuit stability analysis
In impedance acquisition devices, the use of multiplexing chips and the physical characteristics of the PCB inevitably introduce parasitic capacitance at both the inverting input of the operational amplifier in the I-V conversion circuit and the output of the operational amplifier in the voltage excitation circuit. Parasitic capacitance can degrade circuit stability and increase the risk of oscillation. Shen Qidong (1990) studied the impact of parasitic capacitance on operational amplifier circuits and proposed effective compensation methods [27], which serve as valuable references for the circuit analysis in this study.
@@
As shown in Fig.7, the voltage excitation signal $V_{1}$ generates an excitation current through resistor $R_{1}$, which is applied to the inverting input of the operational amplifier in the I-V conversion circuit. Meanwhile, the parasitic capacitance introduced by the multiplexer, which exists between the inverting input of the operational amplifier and ground, influences the circuit’s frequency characteristics. The loop gain of the circuit is:
@@
In the formula, $A_{0}$ and ${\omega_{0}}$ denote the open-loop DC gain and the 3 dB bandwidth of the operational amplifier, respectively. Moreover, $\omega_{P}=\frac{1}{R_{P}\cdot C_{1}}$, $R_{p}=R_{1}\parallel R_{f}$.
In the feedback circuit, the combination of capacitor $C_{1}$ and resistors $R_{1}$ and $R_{f}$ introduces a lagging component. When $C_{1}$ or the feedback resistor $R_{f}$ is relatively large, the second pole frequency associated with the lagging component decreases significantly, making the circuit more susceptible to oscillation. To resolve this, a capacitor $C_{f}$ is typically connected in parallel with the feedback resistor $R_{f}$ for lead compensation. The function of $C_{f}$ is to introduce a zero, and when $C_{f}=\frac{R_{1}\cdot C_{1}}{R_{f}}$, the zero and the pole coincide, thereby compensating the system’s frequency characteristics and effectively preventing circuit instability and oscillation.
@@
As shown in Fig.8, this represents the second-stage section of the voltage excitation circuit. The parasitic capacitance of the multiplexer serves as the capacitive load $C_{L}$ for the operational amplifier, forming a lagging component together with the op-amp’s internal output resistance $R_{0}$. This combination influences the circuit’s frequency characteristics.
@@
Here, ${\omega_{0}}$ represents the intrinsic pole frequency of the operational amplifier, and $\omega_{Q}=\frac{1}{C_{L}\cdot R_{0}}$ denotes the second pole frequency introduced by the output capacitive load and the op-amp’s output resistance. The expression $A_{F}(s)$ is rewritten as:
@@
The damping coefficient $\xi$ of the circuit is expressed as:
@@
where $\omega_{h}$ is: $\omega_{h}=\sqrt{\frac{R_{1}A_{0}}{R_{1}+R_{f}}\omega_{0}\omega_{Q}}$, Since $\omega_{0}\ll\omega_{Q}$, we approximate:
@@
From the expression for the damping coefficient, $\xi$ is inversely proportional to the output resistance $R_{0}$ of the operational amplifier. A lower output resistance $R_{0}$ enhances the damping effect, effectively suppressing system oscillations. The output impedance of the AD8606 increases with both frequency and gain. At 100 kHz with a gain of 100, its output impedance remains below 20 $\Omega$, ensuring compliance with system requirements.

### v2

@@ Abstract
A wireless wearable Electrical Impedance Tomography (EIT) system has been developed utilizing the AD5933 chip to achieve real-time imaging of lung respiration. The system employs a voltage excitation method tailored to human impedance characteristics, injecting current by applying a known voltage and measuring the resulting current through the body. Additionally, specific measures have been implemented to effectively suppress signal oscillations and leakage currents caused by parasitic capacitances. To enhance data acquisition speed, the system employs five parallel AD5933 units, with multiple techniques implemented to ensure high synchronization during simultaneous measurements. Performance testing shows that the system achieves a signal-to-noise ratio greater than 50 dB, a relative standard deviation below 0.3%, and a reciprocity error under 0.8%. Imaging experiments using a water tank phantom, human lungs during breathing, and a resting human calf further demonstrate that this portable EIT system can accurately measure biological tissues with high precision and low cost.
@@ 1 Introduction
EIT is widely used in clinical fields such as lung monitoring [3, 4], brain assessment [5], cancer detection [6], and human-machine interfaces [7]. Among these, lung monitoring is the most common, as EIT provides real-time visualization of pulmonary gas distribution, helping clinicians optimize ventilator settings and reduce the risk of lung injury.
@@
Commercial EIT systems include Sentec LuMon [8], Dräger Pulmovista 500 [9], Timpel Enlight 2100 [10], Gense [11], and Sciospec’s EIT16 and EIT32/64/128+ systems [12]. Notably, the Dräger Pulmovista 500 system provides real-time visualization of pulmonary gas distribution, aiding clinicians in assessing lung collapse, overinflation, and ventilation heterogeneity. Sciospec’s EIT systems typically use 16 electrodes but support configurations up to 256, offering real-time frame rates up to 100 fps and operating across a frequency range of 100 Hz to 1 MHz.
@@
Recent advancements in EIT technology include systems like Sheffield Mk3.5 [13], ACT-4 [14], ACT-5 [15], OXBACT-3 [16], OXBACT-5 [17], UCLH Mk2 [18], UCLH Mk2.5 [19], KHU Mark2 [20], KHU Mark2.5 [21, 22], mfEIT [23], and Dartmouth College-developed systems [24, 25, 26]. Among these, the Sheffield Mk3.5 system can measure impedance at 30 frequency points ranging from 2 kHz to 1.6 MHz, with an average signal-to-noise ratio (SNR) of 40 dB across all frequencies. This system utilizes 8 electrodes and includes 8 identical data acquisition boards, each equipped with a digital signal processor (DSP) to generate driving frequencies and perform fast Fourier transform (FFT) demodulation. The ACT-5 system, equipped with 32 electrodes, can generate conductivity distribution images at a rate of 27 frames per second. Its adaptive current source supports fully programmable current patterns across a frequency range of 5 kHz to 500 kHz, and it can simultaneously display ECG readings during imaging. The KHU Mark2.5 system features automatic calibration and supports continuous operation at a frequency of 250 kHz for up to 24 hours, while maintaining a SNR above 80 dB. Based on FPGA architecture, the system incorporates multiple impedance measurement modules, each equipped with an independent constant current source, a differential voltage measurement unit, and a current source calibrator.
@@
Although the EIT devices mentioned above deliver excellent imaging performance, their high cost limits widespread adoption. Therefore, developing a low-cost, easy-to-implement EIT system is essential. In clinical medicine, EIT is primarily used for real-time monitoring of lung function. Therefore, the device must be compact and portable, with rapid data acquisition capabilities to effectively capture dynamic respiratory changes in the lungs.
@@
This paper aims to develop a portable and low-cost EIT system based on the AD5933 chip. Firstly, according to the impedance characteristics of the human body, current injection is realized by applying an excitation voltage to the object under test and measuring the corresponding excitation current. Compared to traditional methods using a current source, this approach does not require the complete elimination of the DC offset inherent in current sources. Meanwhile, to tackle issues such as oscillations and leakage currents caused by parasitic capacitances in multiplexers, the system implements various optimization measures, enhancing stability at higher frequencies. Secondly, to improve data acquisition speed, the system employs five AD5933 chips operating in parallel, with I2C bus control managed via the TCA9548A and a unified external clock provided by the ICS553, ensuring synchronized data acquisition across all AD5933 channels. Finally, the PCB layout was optimized for a more compact structure. Additionally, Bluetooth wireless communication and lithium battery power were adopted to further improve the device’s portability.
The structure of this paper is organized as follows: Section II introduces the developed EIT system and its specifications. Section III presents experimental results obtained from measurements on a water tank and a male subject. Section IV provides a brief summary.
@@ 2.1 System Architecture
The impedance acquisition system is built around the AD5933, which features an integrated direct digital synthesizer (DDS) capable of generating sine wave signals over a frequency range of 1 kHz to 100 kHz. It also includes a 12-bit analog-to-digital converter (ADC) and a digital signal processor (DSP), allowing for precise measurement of both the amplitude and phase of the sine wave signals.
Fig.1(a) illustrates the overall architecture of the EIT system, while Fig.1(b) shows the physical prototype with all circuit modules clearly labeled. The system consists of an STM32 development board, a power supply module, a Bluetooth transmission module, a voltage excitation module, a buffer module, a data acquisition module, an I2C multiplexer module, an external clock module, and a multiplexer module.
@@
During the impedance acquisition process, the STM32 first uses the I2C bus to control the first AD5933 to generate a sine wave voltage signal. This signal is then converted into an excitation voltage by the voltage excitation module. The output of the voltage excitation circuit and the input of the current measurement circuit are connected to a pair of electrodes via multiplexers to apply voltage excitation. Meanwhile, the AD5933 selects either the output of the voltage measurement circuit or that of the current measurement circuit to acquire the boundary voltage or the current generated by the voltage excitation passing through the object under test. The STM32 controls the channel selection of the multiplexer to perform a full scan of the object under test. The AD5933 transmits the processed voltage and current data to the STM32 via the I2C bus. Finally, the STM32 sends the computed impedance data to a computer via Bluetooth for image reconstruction.
@@
The system is highly flexible and can be configured into different measurement modes based on actual requirements. By adjusting the jumper settings, the enable (EN) pins of the multiplexers connected to the voltage detection circuitry can be set to high or low levels to enable or disable the corresponding multiplexers, thereby reducing the parasitic capacitance caused by unused multiplexers. It can be configured in a two-terminal serial mode for basic impedance analysis, or in a four-terminal serial mode, which separates the current injection and voltage sensing paths to minimize the impact of contact resistance and improve measurement accuracy. Additionally, the system supports a four-terminal parallel mode, enabling simultaneous voltage measurements on multiple electrodes to greatly accelerate data acquisition.
@@
In the two-terminal series measurement mode, impedance is calculated based on the current generated by applying a fixed voltage excitation across the test object. The system architecture for this mode is illustrated in Fig.2.
@@
In the four-terminal series measurement mode, a single voltage excitation allows measurement of either the voltage at one electrode or the current resulting from the voltage excitation passing through the object. Therefore, channel switching via multiplexers is required to measure all boundary voltages and currents under the adjacent excitation configuration. The system architecture for this mode is shown in Fig.3.
@@
In the four-terminal parallel measurement mode, five AD5933 chips are used in parallel, enabling the system to simultaneously acquire the voltages of four electrodes and the current resulting from the voltage excitation passing through the object during a single voltage excitation. By performing multiple channel switches, the system can measure all boundary voltages and currents under the adjacent excitation configuration. The system architecture is shown in Fig.4.
Figure 4: EIT System Architecture in the Four-Terminal Parallel Mode.
@@ 2.2.1 Circuit structure analysis
An excitation voltage is applied to the test object, and the resulting excitation current is measured. In this way, the measured current represents the actual excitation current injected into the test object. Compared to the approach that uses a fixed-amplitude current signal for excitation, this method does not rely on a high output impedance of the current source to stabilize the current amplitude, nor does it require a highly linear relationship between the current amplitude and the control signal to ensure measurement accuracy. Moreover, by employing a fixed-amplitude voltage excitation, the circuit eliminates the need to dynamically adjust the output current in response to the impedance characteristics of the test object. Another benefit of employing voltage excitation is that a small DC bias in the excitation signal results in only a negligible DC current, owing to the human body’s very high impedance to direct current. Consequently, complete elimination of the DC bias voltage is not required. Furthermore, the circuit is both simple and robust, and can operate reliably with a 3.3V power supply.
@@
Fig.5 illustrates the voltage excitation circuit. The output voltage is expressed as follows:
@@
Fig.6 illustrates the I-V conversion circuit. The output voltage is expressed as follows:
@@
The output of the voltage excitation circuit and the current input of the I-V conversion circuit are connected across the test object. The voltage excitation circuit outputs a 1.98 $V_{p-p}$ AC signal superimposed on a DC bias of $\frac{Vcc}{2}$. Since the input of the I-V conversion circuit is also held at $\frac{Vcc}{2}$, the DC voltage across the test object is effectively 0 V, with an AC amplitude of 1.98 $V_{p-p}$.
@@ 2.2.2 Circuit stability analysis
In impedance acquisition devices, the use of multiplexers and the physical characteristics of the PCB inevitably lead to parasitic capacitance at both the input of the I-V conversion circuit and the output of the voltage excitation circuit. Parasitic capacitance can compromise circuit stability and increase the risk of oscillation. Shen Qidong (1990) investigated the effects of parasitic capacitance on operational amplifier circuits and proposed effective compensation methods [27], which provide important references for the circuit analysis in this study.
@@
As shown in Fig.7, the voltage excitation signal $V_{1}$ generates a current through resistor $R_{1}$, which is applied to the input of the I-V conversion circuit. At the same time, the parasitic capacitance from the multiplexer, located between the inverting input of the operational amplifier and ground, affects the circuit’s frequency response. The loop gain of the circuit is:
@@
In the formula, $A_{0}$ and ${\omega_{0}}$ denote the low-frequency open-loop gain and the dominant pole frequency of the operational amplifier, respectively. Moreover, $\omega_{P}=\frac{1}{R_{P}\cdot C_{1}}$, $R_{p}=R_{1}\parallel R_{f}$.
In a feedback circuit, the combination of capacitor $C_{1}$ and resistors $R_{1}$ and $R_{f}$ forms a lagging network. When either $C_{1}$ or the feedback resistor $R_{f}$ is large, the second pole frequency introduced by this network drops significantly, making the circuit more prone to oscillation. To address this issue, a capacitor $C_{f}$ is commonly placed in parallel with the feedback resistor $R_{f}$ to provide lead compensation. $C_{f}$ introduces a zero, and when $C_{f}=\frac{R_{1}\cdot C_{1}}{R_{f}}$, this zero coincides with the pole, effectively compensating the system’s frequency response and helping to prevent instability and oscillation.
@@
As shown in Fig.8, this is the second-stage section of the voltage excitation circuit. The parasitic capacitance from the multiplexer acts as a capacitive load $C_{L}$ on the operational amplifier. Together with the op-amp’s internal output resistance $R_{0}$, it forms a lagging network that affects the circuit’s frequency response.
@@
Here, ${\omega_{0}}$ represents the dominant pole frequency of the operational amplifier, $\omega_{Q}=\frac{1}{C_{L}\cdot R_{0}}$, ${\omega_{Q}}$ denotes the second pole frequency introduced by the capacitive load and the op-amp’s output resistance. The expression $A_{F}(s)$ can be rewritten as:
@@
The damping factor $\xi$ of the circuit is:
@@
where $\omega_{h}=\sqrt{\frac{R_{1}A_{0}}{R_{1}+R_{f}}\omega_{0}\omega_{Q}}$. Since $\omega_{0}\ll\omega_{Q}$, we have:
@@
From the expression of the damping coefficient, it is evident that $\xi$ is inversely proportional to the output resistance $R_{0}$ of the operational amplifier. A lower $R_{0}$ increases the damping effect, thereby helping to suppress system oscillations. The output impedance of the AD8606 rises with both frequency and gain. Even at a frequency of 100 kHz and a gain of 100, the AD8606 maintains an output impedance below 20 $\Omega$, which meets the system’s requirements.